# Road layout in the KPZ class

### v1

@@ 1 Introduction
FPP was originally proposed as a model for the flow of fluid through porous media by Hammersley and Welsh in 1965 [12]. Since then, FPP has evolved into a major area of interest in probability theory. See [1] for an extensive account on significant results in this field. Despite this active interest, many significant unanswered questions remain.
One of the fundamental objects in FPP are geodesics, paths that collect minimal costs connecting two points among a random penalty landscape on the plane. These will be our candidates to model roads. For quantitative analysis of our road model we need refined estimates on geodesics, which unfortunately are not yet available for FPP models. Instead, we switch to exponential last passage percolation (LPP), where extensive results on last passage times and geometry of geodesics are now available. FPP is widely believed to belong to the KPZ universality class – a property proved for LPP. This implies that the geometry of geodesics in the two models should share basic characteristics that we use throughout our arguments. Thus, switching to LPP from FPP seems a reasonable move. Of the vast literature on the KPZ universality class, we refer to the surveys by Corwin [4], Ferrari-Spohn [10], and Quastel [14].
@@
At the end of this paper, we compare our findings to real-world traffic data in Section 6.
@@ 1.1 A Poisson model for road network
This process can be thought of as a marked point process $\xi$ where the starting point $X_{i}$ of car $i$ is a point in the homogeneous Poisson point process on $\mathbb{R}^{2}$, and this gets decorated with the mark $(\theta_{i},\,\ell_{i})$ which represent the direction the car goes towards and the distance it travels respectively. The mark space is $[0,2\pi)\times[0,\infty)$, equipped with the standard product Borel $\sigma$-algebra. By the marking theroem [13, Theorem 5.6], $\xi$ is a Poisson point process on $\bigl{(}\mathbb{R}^{2}\times[0,2\pi)\times[0,\infty)\bigr{)}$. We are after the number $N_{r}$ of cars that come $r$ close to the origin.
@@ 2.1 Further notations and results
$\frac{c_{1}}{n^{1/3}}\leq\mathbb{P}(N\geq n^{4/3})\leq\frac{C_{1}\log n}{n^{1/3}}.$
@@ 3 Bounds for the furthest distance a car can come from
[2, Proposition 1.5] For $\epsilon>0$ there exist $C_{1},c_{1}>0$ such that for all $T>0,\ell>0,n\geq 1$ and for all $\alpha\in(\epsilon,\frac{\pi}{2}-\epsilon)$ we have
@@
[2, Proposition 1.6] For $\psi<1$ there exist $C,c>0$ such that for all $k$ with $|k|+\ell^{1/32}<\psi n^{1/3},$ all $\ell<n^{0.01}$ sufficiently large and all $n\in\mathbb{N}$ sufficiently large we have
@@
Figure 5: To find an upper bound for the second sum in (4.3) we fix $1\leq i<j\leq[\frac{n^{1/3}}{2M}]$. Due to planarity and large transversal fluctuation there is a very small probability that there will be vertices $u$ outside $\mathsf{R}_{i}^{\ell}$ (resp. $\mathsf{R}_{j}^{\ell}$) with $|\phi(v)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ choosing angle in $A_{i}$ (resp. $A_{j}$) such that $\Gamma_{u}^{\theta_{u}}$ will intersect $W$. These are the events $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$. Further, the same argument as in [2, Lemma 3.1] shows that, there is a very small probability that for any geodesic starting from $\mathsf{R}_{i}^{\ell}$(resp. $\mathsf{R}_{j}^{\ell}$) and taking angle in $A_{i}$ (resp. $A_{j}$) will not intersect either $\mathcal{J}_{i}^{\prime}$ or $\mathcal{J}_{i}^{\prime\prime}$ (resp. either $\mathcal{J}_{j}^{\prime}$ or $\mathcal{J}_{j}^{\prime\prime}$). These are the events $\mathcal{G}_{i}$ and $\mathcal{G}_{j}.$ Further, due to Proposition 3.4, we have with very small probability that there are more than $\ell^{1/3}$ distinct vertices in $W$ that has a geodesic starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime}$). Finally, we invoke [2, Lemma 3.2, Lemma 3.3] to conclude that with very small probability any geodesic starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime}$) will have more than $\frac{\ell n}{k_{i,j}^{3}}$ intersection size inside $W$. Combining all the above we get stretched exponential upper bound for each terms in (3.12).
@@
This lemma is already proved in [2, Lemma 3.1], the Reader can refer to this proof. We briefly describe the idea here. We consider two points $a$ and $b$ on $\mathcal{L}_{-n-\lceil\frac{n}{100k_{i,j}}\rceil}$ which are $\lceil 2\ell n^{1/96}n^{2/3}\rceil$ distance away (in the space direction) from the end points of $R_{i}^{\ell}$. Now, by transversal fluctuation estimate we can say that $\Gamma_{a}^{\beta_{i}}$ and $\Gamma_{b}^{\alpha_{i}}$ stays always $\lceil\ell^{1/96}n^{2/3}\rceil$ distance away (in the space direction) from the rectangle $R_{i}^{\ell}$ with probability at least $1-e^{-c\ell^{3/96}}$. On this event, every geodesics starting in $R_{i}^{\ell}$ will be sandwiched between $\Gamma_{a}^{\beta_{i}}$ and $\Gamma_{b}^{\alpha_{i}}$. So, on this event by planarity argument the event $\mathcal{G}_{i}$ will imply large transversal fluctuation of either $\Gamma_{a}^{\beta_{i}}$ or $\Gamma_{b}^{\alpha_{i}}$. This gives us the desired upper bound. ∎
@@
[2, Lemma 3.2, Lemma 3.3] Consider any geodesic $\Gamma_{i}$ (resp. $\Gamma_{j}$) starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime})$. If $I_{\Gamma_{i},\Gamma_{j}}$ denote the intersection size of $\Gamma_{i}$ and $\Gamma_{j}$ inside $W$ then for $\ell$ and $n$ large enough there exist $C,c>0$ such that we have
@@ 4 Bounds for number of cars
In this section we prove the bounds for the tail distribution of $N$.
@@
This proves the upper bound in Theorem 2.4. ∎
@@
Figure 7: To find lower bound for the first sum in (4.3) we fix an $i$. Using Proposition 3.3 we can fix $M$ large enough so that with high probability $\Gamma_{v_{1}}^{\beta_{i}}$ and $\Gamma_{v_{2}}^{\alpha_{i}}$ are contained in $\mathsf{R}_{i}^{M}$ and do not intersect $\mathsf{R}_{i}$. On this event any car starting from $\mathsf{R}_{i}$ and taking angle in $A_{i}$ will intersect $\mathcal{L}_{0}$ on $V_{M}$. Further, using [3, Theorem 3.10] depending on this $M$ we can chose $\ell_{0}$ large enough so that with large probability there are at most $\ell_{0}$ many distinct points on $V_{M}$ carrying a geodesic starting from $V_{M}^{\prime}$ and ending at $V_{M}^{\prime\prime}$. Also, with positive probability there are at least $n^{4/3}$ any vertices $u\in\mathsf{R}_{i}$ such that $\theta_{u}\in A_{i}$. The last event is independent of the first two events. Hence, combining all these we get on a positive probability event there exists $v\in V_{M}$ such that there are at least $\frac{n^{4/3}}{\ell_{0}}$ many $u$ with $\theta_{u}\in A_{i}$ and $v\in\Gamma_{u}^{\theta_{u}}$. Now, applying an averaging argument and summing over all $i$ give a lower bound for the first sum in (4.3).
@@
On $\mathsf{TF}_{i}$, for any $u\in\mathsf{R}_{i}$ with $\theta_{u}\in A_{i},v\in\Gamma_{u}^{\theta_{u}}$, for some $v\in V_{M}.$ Now we define an equivalence class. Let $n_{i}^{\prime\prime}$ denote the intersection point of $y=\tan(\theta_{i})x$ with $\mathcal{L}_{n}$. Consider the two line segments $V_{M}^{\prime}$ (resp. $V_{M}^{\prime\prime}$) with $\lceil 4Mn^{2/3}\rceil$ vertices on $\mathcal{L}_{-n}$ (resp. $\mathcal{L}_{n}$) with midpoint $n_{i}^{\prime}$ (resp. $n_{i}^{\prime\prime}$). For $u,v\in V_{M}^{\prime}$ and $u^{\prime},v^{\prime}\in V_{M}^{\prime\prime},$ we say $(u,v)\sim(u^{\prime},v^{\prime})$ if $\Gamma_{u,v}(0)=\Gamma_{u^{\prime},v^{\prime}}(0).$ Let $E_{i}$ denote the number of equivalence classes. Using [3, Theorem 3.10], we can chose $\ell_{0}$ large enough (depending on $M$) so that for all sufficiently large $n$
@@
As argued before, choosing $M$ sufficiently large completes the proof of Theorem 2.4 lower bound. ∎
@@ 6.1.1 Data acquisition and processing
Our study leverages high-resolution elevation data sourced from the Shuttle Radar Topography Mission (SRTM), accessed through the Earth Explorer platform [15]. The SRTM dataset provides comprehensive global elevation data. From this dataset, we extract elevation data encoded into image files, which serve as input in building an edge weighted graph.
@@ 6.1.2 Graph construction and path analysis
In this study, we leveraged a computational framework found within the GitHub repository [7] to facilitate our analysis of road networks within a designated study area. Utilizing the code from this repository enabled us to construct a weighted graph that models a square lattice segment of the terrain under investigation. Each vertex within this graph represents a precise point on the terrain, and the weights assigned to the edges are calculated based on both the elevation differences and the distances between the vertices at the endpoints.
@@ 6.1.3 Figures
We provide Figures 8, 10 and 10 to illustrate the comparison between actual road segments and theoretical geodesic paths derived from our graph model. Each figure shows a map provided by ESRI World Topo Map [9] with a geodesic path, calculated based on elevation data and shortest path algorithms, overlaid on selected road segments. The intent behind these figures is to visually highlight where and how actual road networks align with or deviate from the calculated geodesic paths. This direct comparison serves as evidence for the hypothesis that road layouts tend to follow geodesic paths, considering elevation.
@@ 6.2 Comparison of the model output and real-world traffic data
We use the government traffic dataset from the United Kingdom accessed at [8]. This consists of annual average daily traffic flow at count points scattered around the roads. We are interested in investigating the validity of Theorems 2.6 and 2.7 for this dataset. We will disregard the logarithmic correction in the former. Fix a startpoint and consider distances $d$ along a spatial direction of LPP. These results say that, with probability separated away from both 0 and 1, we are expected to get at least one road with traffic count proportional to $d^{4}$ within this distance $d$. As opposed to LPP, we consider real road networks to be isotropic, which means that the direction we choose should be irrelevant. We therefore work with East as our preferred direction. The task is to pick points on the map randomly, draw an Easterly line from each, record where this intersects actual roads as well as the traffic volumes carried by these roads.

### v2

# intersection.bib

@@ 1 Introduction
Characteristics of road networks have been considered in numerous works. As in this paper we are interested in a mathematical model to construct road networks, we only mention two papers, first Barthélemy and Flammini [bar_fla_mod_urban_st_patt], who build urban networks by adding new centres and connecting them to the existing network in certain optimal way. Second, Molinero and Hernando [mol_her_model_4_gen_rd] introduce an algorithm based on a succession of Delanuay triangulation of a dense network of the centres of interest plus auxiliary points, and only keeping paths that are shortest between the original nodes of interest in the resulting network.
We build a model where, rather than the locations to connect, the environment surrounding the roads has a crucial effect on the geometry. Points to connect are thought of as having a homogeneous configuration throughout the plane, whereas obstacles roads need to get around will be random which in turn causes nontrivial network characteristics.
@@
FPP was originally proposed as a model for the flow of fluid through porous media by Hammersley and Welsh in 1965 [HW65]. Since then, FPP has evolved into a major area of interest in probability theory. See [ADH15] for an extensive account on significant results in this field. Despite this active interest, many significant unanswered questions remain.
One of the fundamental objects in FPP are geodesics, paths that collect minimal costs connecting two points among a random penalty landscape on the plane. These will be our candidates to model roads. For quantitative analysis of our road model we need refined estimates on geodesics, which unfortunately are not yet available for FPP models. Instead, we switch to exponential last passage percolation (LPP), where extensive results on last passage times and geometry of geodesics are now available. FPP is widely believed to belong to the KPZ universality class – a property proved for LPP. This implies that the geometry of geodesics in the two models should share basic characteristics that we use throughout our arguments. Thus, switching to LPP from FPP seems a reasonable move. Of the vast literature on the KPZ universality class, we refer to the surveys by Corwin [corwin_kpz_univ16], Ferrari-Spohn [fer-spohn_rnd_growth11], and Quastel [quas_intro_kpz12].
LPP models have been subject of intensive research in the last decades. Instead of minimizing weights (which we called costs so far), geodesics are maximising them under the constraint that paths can only take up or right steps. We define this model precisely in Section 2.
The connection between roads and models in the KPZ universality class is not new. Solon, Bunin, Chu and Kardar [sol_bun_chu_kar_opti_paths_poly] compared optimal paths in road networks with directed polymers in random medium (DPRM). LPP can be considered as extremal case of DPRM, and shares similar scaling exponents. Solon et al. find that shortest paths on the road network can be well approximated with DPRM – albeit with the environment exhibiting a power law rather than in exponential LPP. They also conclude that long-range correlations in the road network play an important role in the scaling properties of optimal paths.
@@
The purpose of this paper is to introduce this theoretical traffic model based on LPP, and to rigorously derive some of its interesting characteristics. We do not claim a good fit with reality: in Section 6, we compare our findings to real-world traffic data and point to various reasons why our model stays in the theoretical domain rather than proving useful for practical applications about road networks.
@@ 1.1 A Poisson model for road network
This process can be thought of as a marked point process $\xi$ where the starting point $X_{i}$ of car $i$ is a point in the homogeneous Poisson point process on $\mathbb{R}^{2}$, and this gets decorated with the mark $(\theta_{i},\,\ell_{i})$ which represent the direction the car goes towards and the distance it travels respectively. The mark space is $[0,2\pi)\times[0,\infty)$, equipped with the standard product Borel $\sigma$-algebra. By the marking theroem [last2017lectures, Theorem 5.6], $\xi$ is a Poisson point process on $\bigl{(}\mathbb{R}^{2}\times[0,2\pi)\times[0,\infty)\bigr{)}$. We are after the number $N_{r}$ of cars that come $r$ close to the origin.
@@ 2.1 Further notations and results
$\frac{c_{1}}{n^{1/4}}\leq\mathbb{P}(N\geq n)\leq\frac{C_{1}\log n}{n^{1/4}}.$
@@ 3 Bounds for the furthest distance a car can come from
[BBB23, Proposition 1.5] For $\epsilon>0$ there exist $C_{1},c_{1}>0$ such that for all $T>0,\ell>0,n\geq 1$ and for all $\alpha\in(\epsilon,\frac{\pi}{2}-\epsilon)$ we have
@@
[BBB23, Proposition 1.6] For $\psi<1$ there exist $C,c>0$ such that for all $k$ with $|k|+\ell^{1/32}<\psi n^{1/3},$ all $\ell<n^{0.01}$ sufficiently large and all $n\in\mathbb{N}$ sufficiently large we have
@@
Figure 5: To find an upper bound for the second sum in (4.3) we fix $1\leq i<j\leq[\frac{n^{1/3}}{2M}]$. Due to planarity and large transversal fluctuation there is a very small probability that there will be vertices $u$ outside $\mathsf{R}_{i}^{\ell}$ (resp. $\mathsf{R}_{j}^{\ell}$) with $|\phi(v)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil$ choosing angle in $A_{i}$ (resp. $A_{j}$) such that $\Gamma_{u}^{\theta_{u}}$ will intersect $W$. These are the events $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$. Further, the same argument as in [BBB23, Lemma 3.1] shows that, there is a very small probability that for any geodesic starting from $\mathsf{R}_{i}^{\ell}$(resp. $\mathsf{R}_{j}^{\ell}$) and taking angle in $A_{i}$ (resp. $A_{j}$) will not intersect either $\mathcal{J}_{i}^{\prime}$ or $\mathcal{J}_{i}^{\prime\prime}$ (resp. either $\mathcal{J}_{j}^{\prime}$ or $\mathcal{J}_{j}^{\prime\prime}$). These are the events $\mathcal{G}_{i}$ and $\mathcal{G}_{j}.$ Further, due to Proposition 3.4, we have with very small probability that there are more than $\ell^{1/3}$ distinct vertices in $W$ that has a geodesic starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime}$). Finally, we invoke [BBB23, Lemma 3.2, Lemma 3.3] to conclude that with very small probability any geodesic starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime}$) will have more than $\frac{\ell n}{k_{i,j}^{3}}$ intersection size inside $W$. Combining all the above we get stretched exponential upper bound for each terms in (3.12).
@@
This lemma is already proved in [BBB23, Lemma 3.1], the Reader can refer to this proof. We briefly describe the idea here. We consider two points $a$ and $b$ on $\mathcal{L}_{-n-\lceil\frac{n}{100k_{i,j}}\rceil}$ which are $\lceil 2\ell n^{1/96}n^{2/3}\rceil$ distance away (in the space direction) from the end points of $R_{i}^{\ell}$. Now, by transversal fluctuation estimate we can say that $\Gamma_{a}^{\beta_{i}}$ and $\Gamma_{b}^{\alpha_{i}}$ stays always $\lceil\ell^{1/96}n^{2/3}\rceil$ distance away (in the space direction) from the rectangle $R_{i}^{\ell}$ with probability at least $1-e^{-c\ell^{3/96}}$. On this event, every geodesics starting in $R_{i}^{\ell}$ will be sandwiched between $\Gamma_{a}^{\beta_{i}}$ and $\Gamma_{b}^{\alpha_{i}}$. So, on this event by planarity argument the event $\mathcal{G}_{i}$ will imply large transversal fluctuation of either $\Gamma_{a}^{\beta_{i}}$ or $\Gamma_{b}^{\alpha_{i}}$. This gives us the desired upper bound. ∎
@@
[BBB23, Lemma 3.2, Lemma 3.3] Consider any geodesic $\Gamma_{i}$ (resp. $\Gamma_{j}$) starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime})$. If $I_{\Gamma_{i},\Gamma_{j}}$ denote the intersection size of $\Gamma_{i}$ and $\Gamma_{j}$ inside $W$ then for $\ell$ and $n$ large enough there exist $C,c>0$ such that we have
@@ 4 Bounds for number of cars
In this section we prove the bounds for the tail distribution of $N$. In particular, we will prove there exist $c_{1},C_{1}>0$ (depending on $\varepsilon$) such that for all $n$ sufficiently large
$\frac{c_{1}}{n^{1/3}}\leq\mathbb{P}\left(N\geq n^{4/3}\right)\leq\frac{C_{1}\log n}{n^{1/3}}.$
Note that, this will imply Theorem 2.4.
@@
This proves the upper bound in Theorem 2.4 after an $n\leftrightarrow n^{4/3}$ change of variable. ∎
@@
Figure 7: To find lower bound for the first sum in (4.3) we fix an $i$. Using Proposition 3.3 we can fix $M$ large enough so that with high probability $\Gamma_{v_{1}}^{\beta_{i}}$ and $\Gamma_{v_{2}}^{\alpha_{i}}$ are contained in $\mathsf{R}_{i}^{M}$ and do not intersect $\mathsf{R}_{i}$. On this event any car starting from $\mathsf{R}_{i}$ and taking angle in $A_{i}$ will intersect $\mathcal{L}_{0}$ on $V_{M}$. Further, using [BHS18, Theorem 3.10] depending on this $M$ we can chose $\ell_{0}$ large enough so that with large probability there are at most $\ell_{0}$ many distinct points on $V_{M}$ carrying a geodesic starting from $V_{M}^{\prime}$ and ending at $V_{M}^{\prime\prime}$. Also, with positive probability there are at least $n^{4/3}$ any vertices $u\in\mathsf{R}_{i}$ such that $\theta_{u}\in A_{i}$. The last event is independent of the first two events. Hence, combining all these we get on a positive probability event there exists $v\in V_{M}$ such that there are at least $\frac{n^{4/3}}{\ell_{0}}$ many $u$ with $\theta_{u}\in A_{i}$ and $v\in\Gamma_{u}^{\theta_{u}}$. Now, applying an averaging argument and summing over all $i$ give a lower bound for the first sum in (4.3).
@@
On $\mathsf{TF}_{i}$, for any $u\in\mathsf{R}_{i}$ with $\theta_{u}\in A_{i},v\in\Gamma_{u}^{\theta_{u}}$, for some $v\in V_{M}.$ Now we define an equivalence class. Let $n_{i}^{\prime\prime}$ denote the intersection point of $y=\tan(\theta_{i})x$ with $\mathcal{L}_{n}$. Consider the two line segments $V_{M}^{\prime}$ (resp. $V_{M}^{\prime\prime}$) with $\lceil 4Mn^{2/3}\rceil$ vertices on $\mathcal{L}_{-n}$ (resp. $\mathcal{L}_{n}$) with midpoint $n_{i}^{\prime}$ (resp. $n_{i}^{\prime\prime}$). For $u,v\in V_{M}^{\prime}$ and $u^{\prime},v^{\prime}\in V_{M}^{\prime\prime},$ we say $(u,v)\sim(u^{\prime},v^{\prime})$ if $\Gamma_{u,v}(0)=\Gamma_{u^{\prime},v^{\prime}}(0).$ Let $E_{i}$ denote the number of equivalence classes. Using [BHS18, Theorem 3.10], we can chose $\ell_{0}$ large enough (depending on $M$) so that for all sufficiently large $n$
@@
As argued before, choosing $M$ sufficiently large completes the proof of Theorem 2.4 lower bound (with $n\leftrightarrow n^{4/3}$, as before). ∎
@@ 6.1.1 Data acquisition and processing
Our study leverages high-resolution elevation data sourced from the Shuttle Radar Topography Mission (SRTM), accessed through the Earth Explorer platform [earthexplorer]. The SRTM dataset provides comprehensive global elevation data. From this dataset, we extract elevation data encoded into image files, which serve as input in building an edge weighted graph.
@@ 6.1.2 Graph construction and path analysis
In this study, we leveraged a computational framework found within the GitHub repository [statmechmodels] to facilitate our analysis of road networks within a designated study area. Utilizing the code from this repository enabled us to construct a weighted graph that models a square lattice segment of the terrain under investigation. Each vertex within this graph represents a precise point on the terrain, and the weights assigned to the edges are calculated based on both the elevation differences and the distances between the vertices at the endpoints.
@@ 6.1.3 Figures
We provide Figures 8, 10 and 10 to illustrate the comparison between actual road segments and theoretical geodesic paths derived from our graph model. Each figure shows a map provided by ESRI World Topo Map [esriworldtopo2024] with a geodesic path, calculated based on elevation data and shortest path algorithms, overlaid on selected road segments. The intent behind these figures is to visually highlight where and how actual road networks align with or deviate from the calculated geodesic paths. This direct comparison serves as evidence for the hypothesis that road layouts tend to follow geodesic paths, considering elevation.
@@ 6.2 Comparison of the model output and real-world traffic data
We use the government traffic dataset from the United Kingdom accessed at [roadtraffic_gov_uk]. This consists of annual average daily traffic flow at count points scattered around the roads. We are interested in investigating the validity of Theorems 2.6 and 2.7 for this dataset. We will disregard the logarithmic correction in the former. Fix a startpoint and consider distances $d$ along a spatial direction of LPP. These results say that, with probability separated away from both 0 and 1, we are expected to get at least one road with traffic count proportional to $d^{4}$ within this distance $d$. As opposed to LPP, we consider real road networks to be isotropic, which means that the direction we choose should be irrelevant. We therefore work with East as our preferred direction. The task is to pick points on the map randomly, draw an Easterly line from each, record where this intersects actual roads as well as the traffic volumes carried by these roads.